### v1

@@ 5.4 Onset of Sickness & Correlations
The FMS scores increased throughout sickness stimulation, but the effect appeared to diminish toward the end of the exposure phase, suggesting a non-linear trajectory (see Fig. 3). Physiological measures followed the general trend of a reaction: HR increased during stimulation, while HRV decreased, both returning to baseline levels during the recovery phase. EDA also increased with stimulation but exhibited a more delayed reaction, remaining elevated throughout the recovery period. This general pattern is consistent with prior findings in real-world settings and VR research [10, 6, 50].
@@
Concerning endocrinological response (salivary cortisol), the only significant predictor in all regression models was the variable coding voluntary termination of VR motion by the participant. No other predictors—including HR, HRV, or EDA—were statistically significant. This outcome is not necessarily unexpected and aligns with earlier findings, especially when considering the relatively small sample size (), which limits the detection of subtle effects such as those associated with sex differences [27, 23].
@@ 5.5 Limitations
In addition to the limitations already discussed, our study design presents several constraints that should be considered.
A low-level (passive) control condition was chosen to account for the well-known circadian rhythm of the HPA axis [35]. Consequently, it cannot be ruled out that part of the observed response was influenced by unspecific contextual factors of the test situation, such as the VR environment or the presence of the experimenter. Future studies should incorporate active control conditions to better quantify the specific effects of HPA activation induced by rotational movement in VR.
Furthermore, our stimulation lasted a maximum of eight minutes and was not designed to push participants to their individual limits. This implies that participants may not have experienced the full spectrum of cybersickness symptoms. A meta-analysis by [59] categorized studies into exposure durations of 0–10 minutes, 10–20 minutes, and beyond 20 minutes. Their findings indicated that while oculomotor symptoms remained stable across these groups, nausea and disorientation peaked between 10 and 20 minutes. This suggests that our results may not be fully generalizable to longer exposure durations or more intense sickness conditions.

### v2

@@ 5.4 Onset of Sickness & Correlations
The FMS scores increased throughout sickness stimulation, but the effect appeared to diminish toward the end of the exposure phase, suggesting a non-linear trajectory (see Fig. 3). Physiological measures followed the general trend of a reaction: HR increased during stimulation, while HRV decreased, both returning to baseline levels during the recovery phase. EDA also increased with stimulation but exhibited a more delayed reaction, remaining elevated throughout the recovery period. This general pattern is consistent with prior findings in real-world settings and VR research [Dennison2016, Bos2005motion, Munafo2017TheVR].
@@
Concerning endocrinological response (salivary cortisol), the only significant predictor in all regression models was the variable coding voluntary termination of VR motion by the participant. No other predictors—including HR, HRV, or EDA—were statistically significant. This outcome is not necessarily unexpected and aligns with earlier findings, especially when considering the relatively small sample size (), which limits the detection of subtle effects such as those associated with sex differences [Kelly2023gender, Howard2021].
@@ 5.5 Limitations
In addition to the limitations already discussed, our study design presents several constraints that should be considered. A low-level (passive), time-matched control condition was chosen to obtain a conservative baseline for endocrine measures that are highly sensitive to contextual arousal and environmental stimulation. However, this design cannot fully disentangle cybersickness-specific effects from other unspecific aspects of the test situation (e.g., VR context, experimenter presence). Future studies should therefore include an additional active VR control condition to separate VR-related activation from cybersickness-specific effects. In addition, the exploratory design and moderate sample size limit the generalizability of effect size estimates, particularly given substantial interindividual differences in cybersickness susceptibility and recent recommendations to aim for higher statistical power in cybersickness research [Weech2019].
Furthermore, our stimulation lasted a maximum of eight minutes and was not designed to push participants to their individual limits. This implies that participants may not have experienced the full spectrum of cybersickness symptoms. A meta-analysis by [Saredakis2020factors] categorized studies into exposure durations of 0–10 minutes, 10–20 minutes, and beyond 20 minutes. Their findings indicated that while oculomotor symptoms remained stable across these groups, nausea and disorientation peaked between 10 and 20 minutes. This suggests that our results may not be fully generalizable to longer exposure durations or more intense sickness conditions.